OpenStudy (anonymous):

A child is creating a pyramid with building blocks. The top three levels include 5 blocks, 12 blocks, and 19 blocks. How many blocks would be needed for a pyramid 25 levels tall? Use complete sentences to explain how a sum of an arithmetic series was applied.

OpenStudy (anonymous):

[5 + (7*0)] + [5 + (7*1)] + [5 + (7*2)] + ... + [5 + (7*24)] = (25)(5) + 7(1 + 2 + 3 + ... + 24) = 125 + 7[(24)(25)]/2

OpenStudy (anonymous):

That's a lot of blocks for one child! :-)

OpenStudy (anonymous):

The key is that the sum of 1 + 2 + 3 + ... + (n-1) + n = n(n+1)/2

OpenStudy (anonymous):

Making sense? If not, ask any question you like.